On January 1, 2021, the company obtained a $3 million loan with a 12% interest rate. The building was completed on SeptemberRequired: 1. Calculate the amount of interest that Mason should capitalize in 2021 and 2022 using the weighted-average methodReq 1 and 3 Req 2 What is the total cost of the building? (Round Weighted-average rate of all debt to 2 decimal places but
Answer #1

Capitalization period C=B/12 Weighted Average Accumulated Expenditures D = (A*C) Year 2021 Calculation of weighted Average AcIn Year 2022 Beginning Expenditure cost (3,000,000 + 192, 252) = $ 3,192,252 Calculation of weighted Average Accumulated Expe1,200,000 690,000 450,000 660,000 2) Cost of Building Expenditures incurred in 2021 January 01, 2021 March 01, 2021 June 30,